# Polycubes via Dual Loops

### v1

@@ 3 Polycube dual characterization
      For the sake of contradiction, assume that two edges $(v,u_{1})$ and $(v,u_{2})$ incident on the same corner $v$ overlap, and assume without loss of generality that they both point in the positive $X$-direction. Then both $u_{1}$ and $u_{2}$ must have been assigned a higher $x$-coordinate than $v$, and hence the $X$-graph must contain a directed edge from the zone of $v$ to the zone of $u_{1}$, and also from the zone of $v$ to the zone of $u_{2}$. But then the loop region of $\mathcal{L}$ corresponding to $v$ must have two loop segments on the boundary with label $X$ and side --. This contradicts Condition 3 of Definition 6.
    Thus, $Q$ is indeed a polycube corresponding to $\mathcal{L}$. Now assume that two polycubes $Q_{1}$ and $Q_{2}$ correspond to $\mathcal{L}$. Then the quad meshes of $Q_{1}$ and $Q_{2}$ must be isomorphic, and the positions of the corners must adhere to the (strict) partial orders induced by the $X$-, $Y$-, and $Z$-graphs of $\mathcal{L}$. But then $Q_{1}$ and $Q_{2}$ must be order-equivalent by Definition 4.
@@
Theorem 11
For any polycube loop structure $\mathcal{L}$, the $X$-, $Y$-, and $Z$-graphs each satisfy $|E|=g+|V|-1$, where $g$ is the genus of the corresponding polycube $Q$, and $|E|$ and $|V|$ indicate the number of edges and vertices in a graph, respectively.
@@
    If $\mathcal{S}$ is decomposed into $k$ subsurfaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{k}$ by cutting along a collection of non-intersecting loops, where a subsurface $\mathcal{S}_{i}$ has Euler characteristic $\chi_{i}=2-2g_{i}-b_{i}$, then we can obtain $\chi$ using the additive form:
@@
    $\chi=2n-2l+\sum^{k}_{i=1}(-2g_{i})$
    Furthermore, the zones in a polycube loop structure have genus 0 by definition, and hence we know that $\sum^{k}_{i=1}(-2g_{i})=0$. Thus we obtain that $\chi=2k-2l$. Given that $\chi=2-2g$, we derive the relation $2-2g=2k-2l$, or equivalently, $l=g+k-1$. The $X$-, $Y$-, and $Z$-graphs have $|V|=k$ and $|E|=l$. Therefore, we know that for the $X$-, $Y$-, and $Z$-graph each satisfy $|E|=g+|V|-1$.
Theorem 11 also directly implies the following nice property for polycubes of genus 00.
Corollary 12
@@
The conditions of Definition 6 that are affected by the orientation of the loops are Conditions 3 and 6. Note that we can still compute the $X$-, $Y$-, $Z$-graphs of a labeled loop structure, but we cannot assign directions to the edges in the graphs. To assign orientations to the loops, we construct a graph on the loops for each axis separately ($X$, $Y$, or $Z$). For the $X$-axis, we make a graph that has two nodes for each $X$-loop, one corresponding to each side of the loop. We connect the two nodes of a loop with an edge, and we also connect two nodes if the corresponding loop sides both occur on the boundary of a single loop region. If this graph admits a $2$-coloring (that is, if it is bipartite) for all axes, then we can consistently assign side labels ($+$/--) to the sides of loops, or equivalently, orient the loops such that Condition 3 of Definition 6 is always satisfied. Finally, we must also check if the $X$-, $Y$-, $Z$-graphs are acyclic. Note that, by Corollary 12, this condition can never be violated for polycubes of genus 00. For polycubes of higher genus, the challenge is to find a $2$-coloring of each of the $3$ graphs described above such that the $X$-, $Y$-, $Z$-graphs become acyclic. If the graph to be colored has $k$ connected components, then the graph has $2^{k}$ possible $2$-colorings. So, even though we can exhaustively check if a labeled loop structure can be oriented to become a polycube loop structure, it remains open if this can be computed efficiently.
Figure 9: Sequence of loop additions. A $Z$-loop is added, then a $Y$-loop, and then a $X$-loop.
Polycube corners. For completeness we show all possible polycube corner configurations that can occur and how they are represented in a polycube loop structure. At each polycube corner, the edges align with one of the coordinate axes ($X$, $Y$, or $Z$) and extend in either the positive or negative direction.
@@
We can classify each configuration in six general patterns (see Figure 8). The simple corner with only 3 adjacent edges; The edge corner and flat corner with 4 adjacent edges (one using all three axes, the other using only two axes); The bend corner with 5 adjacent edges; and the symmetric and asymmetric complex corner (one with a symmetric pattern, other with an asymmetric pattern).
@@ 4 Polycube modifications
To construct a polycube map or polycube segmentation for an input surface $\mathcal{S}$, we have to choose a suitable polycube $Q$ to map from. Many current methods for constructing polycube segmentations do not keep track of the corresponding polycube $Q$ while they are constructing the segmentation. This creates friction between optimizing the quality of the polycube segmentation (complexity and alignment) and guaranteeing the validity of the polycube segmentation (that it actually corresponds to a valid polycube). In fact, many methods ignore the validity of the polycube segmentation until it has been fully optimized, and require an additional fixing step if the segmentation turns out to be invalid, which may have a significant negative effect on the final quality.
Instead, we suggest to construct the polycube segmentation iteratively, where we guarantee the validity of the segmentation during every step. Doing this directly using the polycube segmentation on $\mathcal{S}$ is challenging, but our dual characterization of polycubes offers a relatively straightforward approach: we initialize a simple polycube loop structure $\mathcal{L}$ (with the same genus as $\mathcal{S}$) on the surface $\mathcal{S}$, and then we iteratively add or remove oriented and labeled loops to $\mathcal{L}$ whilst ensuring that it remains a polycube loop structure. We can then also easily keep track of the corresponding polycube $Q$ using Theorem 10. The loops we add or remove can be chosen based on some quality criteria for the final polycube segmentation. Although this approach yields only a polycube loop structure $\mathcal{L}$ (on $\mathcal{S}$), this loop structure prescribes the global structure of the polycube segmentation and only local optimizations are needed to obtain the final polycube segmentation.
Although we can simply check the full characterization of polycube loop structures after every step, this is expensive, as we will only know after the check whether the addition or removal was allowed. This is especially problematic for adding loops, as there are many potential options for adding a loop. Therefore we show how to effectively construct loops that can be added whilst keeping $\mathcal{L}$ valid and how to check efficiently if a loop can be removed from $\mathcal{L}$.
@@
The addition of a loop to a polycube loop structure $\mathcal{L}$ introduces a strip of faces in the corresponding polycube $Q$, see Figure 9. We say that a loop is valid if it can be added to $\mathcal{L}$ such that $\mathcal{L}$ remains a polycube loop structure (according to Definition 6). To effectively find valid loops for $\mathcal{L}$, we develop a directed graph structure that enables the enumeration of all possible valid loops.
We say that any two loops are combinatorially equivalent if their intersection pattern (the loop segments they intersect) in $\mathcal{L}$ are equivalent. As such, we observe that all possible (valid or invalid) loops that could be added to a loop structure are characterized by the edge-graph of $\mathcal{L}$, which is defined as follows: there is a vertex for each loop segment in $\mathcal{L}$, and there is a directed edge from vertex $u$ to vertex $v$ if the two corresponding loop segments bound the same loop region in $\mathcal{L}$. We call this directed graph $G_{L}$. Note that this graph is directed since we also care about the orientation of the loops. For an oriented loop, we assume that the right side (as seen from the direction of the loop) is the positive ($+$) side, and the left side is the negative (--) side.
The validity of a loop in $G_{L}$ also depends on its label. Therefore, we will construct three different graphs, $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$, one for each possible label, where each graph is obtained by removing directed edges from $G_{L}$. The idea is that a loop with label $\Delta\in\{X,Y,Z\}$ is valid if and only if it corresponds to a simple (oriented) cycle in $G_{V}^{\Delta}$, where we define a cycle to be simple if it visits every loop region of $\mathcal{L}$ at most once.
@@
Local constraints. When we add a loop $\lambda$ to $\mathcal{L}$, all loop regions that are visited will be split into two loop regions. All other loop regions are unaffected. Furthermore, the single zone that encompasses the loop will be split into two zones. Note that splitting a surface of genus 00 cannot result in regions of higher genus, so Conditions 4 and 5 of Definition 6 are never violated when adding a loop.
The other local conditions of Definition 6 (1-3) can be guaranteed by constructing the correct graphs $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$. First of all, Condition 1 can always be guaranteed by construction, as there always exists a combinatorially equivalent loop with that property. For Conditions 2 and 3, observe that every (directed) edge in $G_{L}$ corresponds to an individual split of a loop region into two loop regions. For each label $\Delta\in\{X,Y,Z\}$ of the loop we can simply directly check, for every directed edge in $G_{L}$, if the corresponding split results in valid loop regions according to Conditions 2 and 3. We include a directed edge in $G_{V}^{\Delta}$ if and only if the corresponding split is valid. There are three cases to consider based on the number of $\Delta$-loop segments within the loop region. If the loop region contains two $\Delta$-loop segments, the edge must separate the two existing loops. If the loop region contains one $\Delta$-loop segment, the edge must not intersect the existing $\Delta$-loop segment. If the loop region contains no $\Delta$-loop segments, there are no conflicts, and all edges are permitted. There is one consistent orientation of the edges if the loop region contains one or two $\Delta$-loop segments, such that the new edge does not conflict with existing edges. Otherwise, edges can be directed in either direction. These three general cases are illustrated in Figure 10, where all valid outgoing edges for a single loop segment (vertex in graph $G_{V}^{X}$) are shown.
Note that we need to require that every loop region is visited at most once; otherwise, there will be a new loop region where the same loop (the added loop) occurs twice on the boundary. Since the underlying surface $\mathcal{S}$ is orientable, the side label of the two loop segments in the direction of that loop region must also be the same, which would violate Condition 3 of Definition 6. Thus, a loop with label $\Delta\in\{X,Y,Z\}$ satisfies Conditions 1-5 of Definition 6 if and only if it corresponds to a directed cycle in $G_{V}^{\Delta}$ that visits every loop region at most once (simple).
Maintaining acyclicity. The only condition that remains to be checked is Condition 6 of Definition 6. However, we show that adding a loop can never introduce a cycle in the $X$-, $Y$-, or $Z$-graphs. Hence, a loop with label $\Delta\in\{X,Y,Z\}$ is valid if and only if it corresponds to a non-trivial simple directed cycle in $G_{V}^{\Delta}$.
@@
First of all, Condition 1 of Definition 5 cannot be violated by the removal of a loop. For the other local Conditions 2-4, observe that, when removing a loop $\lambda$ from $\mathcal{L}$, every loop segment of $\lambda$ causes two loop regions to be merged. We can easily find these loop regions efficiently by tracing $\lambda$ through $\mathcal{L}$ and then we can simply check if the resulting loop regions satisfy Conditions 2-4. Note that Condition 4 can never be violated by removing $\lambda$, since otherwise there must be a loop segment of $\lambda$ for which the same loop region is on both sides. This would correspond to a corner of the corresponding polycube $Q$ connected by an edge to itself, which is clearly not possible.
Condition 5 of Definition 6 can also not be violated by removing a loop $\lambda$ from $\mathcal{L}$. Assume for the sake of contradiction that there is an $X$-zone $\zeta$ after removing $\lambda$ with higher genus. If $\zeta$ had genus 00 before removing $\lambda$, then, by the definition of genus, $\lambda$ must cut $\zeta$ in such a way that $\zeta$ remains connected. But then, by the construction of the $X$-graph, the vertex $v$ corresponding to $\zeta$ must have a self-loop caused by $\lambda$ before removing $\lambda$. This contradicts Condition 6 of Definition 6.
Finally, we need to check Condition 6 of Definition 6 after removing $\lambda$ from $\mathcal{L}$. Note that $\lambda$ corresponds to an edge in the $X$-graph. Removing $\lambda$ does not only remove this edge, but also contracts the edge (since the zones corresponding to the endpoints are merged) into a new vertex $v$. This may introduce a cycle in the $X$-graph through $v$, but this can efficiently be checked using a simple depth first search from $v$.
@@
We first consider polycubes of genus 00, where the initial polycube loop structure corresponds to a single cube. In Figure 11(a) we show a polycube that cannot be obtained from the single cube by only adding loops. We refer to this polycube as the non-incremental polycube. It can easily be verified that none of the loops of the non-incremental polycube can be removed without violating the conditions of a polycube loop structure. However, we can construct the non-incremental polycube by both adding and removing loops. In fact, we do not know if all polycubes of genus 00 can be obtained by adding and removing loops, starting from the single cube. We leave this as an open problem.
For polycubes of higher genus, the holes of the polycube play an important role. Specifically, the orientation of a hole chosen in the initial polycube cannot be changed by adding or removing loops (see Figure 11(b, c)). Therefore, for polycube segmentation of surfaces with higher genus, it is important to pick the right orientation of each hole already in the initial polycube. Alternatively, we can introduce other operations on polycube loop structures that can introduce/eliminate a hole or change its orientation, but we leave this as future work.
Figure 12: Polycube segmentations constructed with our proof-of-concept implementation. We show the embedded polycube loop structure, the corresponding polycube, and the final polycube segmentation.

### v3

@@ 3 Polycube dual characterization
      For the sake of contradiction, assume that two edges $(v,u_{1})$ and $(v,u_{2})$ incident on the same corner $v$ overlap, and assume without loss of generality that they both point in the positive $X$-direction. Then both $u_{1}$ and $u_{2}$ must have been assigned a higher $x$-coordinate than $v$, and hence the $X$-graph must contain a directed edge from the zone of $v$ to the zone of $u_{1}$, and also from the zone of $v$ to the zone of $u_{2}$. But then the loop region of $\mathcal{L}$ corresponding to $v$ must have two loop segments on the boundary with label $X$ and side --. This contradicts Condition 3 of Definition 3.2.
    Thus, $Q$ is indeed a polycube corresponding to $\mathcal{L}$. Now assume that two polycubes $Q_{1}$ and $Q_{2}$ correspond to $\mathcal{L}$. Then the quad meshes of $Q_{1}$ and $Q_{2}$ must be isomorphic, and the positions of the corners must adhere to the (strict) partial orders induced by the level graphs of $\mathcal{L}$. But then $Q_{1}$ and $Q_{2}$ must be order-equivalent by Definition 2.2.
@@
Theorem 3.4
For any polycube loop structure $\mathcal{L}$, the $X$-, $Y$-, and $Z$-graphs each satisfy $|E|=g+|V|-1-\Gamma$, where $|E|$ and $|V|$ indicate the number of edges and vertices in a graph, $g$ is the genus of the corresponding polycube $Q$, and $\Gamma$ is the sum of genii for each zone.
@@
    If $\mathcal{S}$ is decomposed into $k$ subsurfaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{k}$ by cutting along a collection of non-intersecting loops, where a subsurface $\mathcal{S}_{i}$ has Euler characteristic: $\chi_{i}=2-2g_{i}-b_{i}$.
    Then we can obtain $\chi$ using the additive form:
@@
    $\chi=2k-2l+\sum^{k}_{i=1}(-2g_{i})$
    Given that $\chi=2-2g$, we derive the relation $2-2g=2k-2l+\sum^{k}_{i=1}(-2g_{i})$, or equivalently, $l=g+k-1-\sum^{k}_{i=1}g_{i}$. The $X$-, $Y$-, and $Z$-graphs have $|V|=k$ and $|E|=l$. Recall $\Gamma$ is defined as $\sum^{k}_{i=1}g_{i}$. Therefore, we know that for the $X$-, $Y$-, and $Z$-graph each satisfy $|E|=g+|V|-1-\Gamma$.
Theorem 3.4 also directly implies the following nice property for polycubes of genus 00.
Corollary 3.1
@@
Figure 9: Sequence of loop additions: a $Z$-loop, then a $Y$-loop, and then a $X$-loop. Note that loops can be added between existing loops, as in the second figure.
@@
The conditions of Definition 3.2 that are affected by the orientation of the loops are Conditions 3 and 5. Note that we can still compute the $X$-, $Y$-, $Z$-graphs of a labeled loop structure, but we cannot assign directions to the edges in the graphs. To assign orientations to the loops, we construct a graph on the loops for each axis separately ($X$, $Y$, or $Z$). For the $X$-axis, we make a graph that has two nodes for each $X$-loop, one corresponding to each side of the loop. We connect the two nodes of a loop with an edge, and we also connect two nodes if the corresponding loop sides both occur on the boundary of a single loop region. If this graph admits a $2$-coloring (that is, if it is bipartite) for all axes, then we can consistently assign side labels ($+$/--) to the sides of loops, or equivalently, orient the loops such that Condition 3 of Definition 3.2 is always satisfied. Finally, we must also check Condition 5 of Definition 3.2: the $X$-, $Y$-, $Z$-graphs are acyclic. Note that, by Corollary 3.1, this condition can never be violated for polycubes of genus 00. For polycubes of higher genus, the challenge is to find a $2$-coloring of each of the $3$ graphs described above such that the $X$-, $Y$-, $Z$-graphs become acyclic. If the graph to be colored has $k$ connected components, then the graph has $2^{k}$ possible $2$-colorings. So, even though we can exhaustively check if a labeled loop structure can be oriented to become a polycube loop structure, it remains open if this can be computed efficiently.
Polycube corners. For completeness we show all possible polycube corner configurations that occur and how they are represented in a polycube loop structure. At each polycube corner, the edges align with one of the coordinate axes ($X$, $Y$, or $Z$) and extend in either the positive or negative direction.
@@
We can classify each configuration in six general patterns (see Figure 8). The simple corner with only 3 adjacent edges; The edge corner and flat corner with 4 adjacent edges (one using all three axes, the other using only two axes); The bent corner with 5 adjacent edges; and the symmetric and asymmetric complex corner (one with a symmetric intersection pattern, the other with an asymmetric intersection pattern).
@@ 4 Polycube modifications
To construct a polycube map or polycube segmentation for an input surface $\mathcal{S}$, we have to choose a suitable polycube $Q$ to map from. Current methods for constructing polycube segmentations do not have a guarantee on the existence of a corresponding polycube $Q$ while they are constructing the segmentation. This creates friction between optimizing the quality of the polycube segmentation (complexity and alignment) and guaranteeing the validity of the polycube segmentation (that it actually corresponds to a valid polycube).
Instead, we suggest to construct the polycube segmentation iteratively, where we guarantee the validity of the segmentation during every step. Doing this directly using the polycube segmentation on $\mathcal{S}$ is challenging, but our dual characterization of polycubes offers a relatively straightforward approach: we initialize a simple polycube loop structure $\mathcal{L}$ (with the same genus as $\mathcal{S}$) on the surface $\mathcal{S}$, and then we iteratively add or remove oriented and labeled loops to $\mathcal{L}$ whilst ensuring that it remains a polycube loop structure. We can then also easily keep track of the corresponding polycube $Q$ using Theorem 3.3. The loops we add or remove can be chosen based on some quality criteria for the final polycube segmentation. Although this approach yields only a polycube loop structure $\mathcal{L}$ (on $\mathcal{S}$), this loop structure prescribes the global structure of the polycube segmentation and only local optimizations are needed to obtain the final polycube segmentation.
We can simply check the characterization after each step to determine if an addition or removal is valid. Doing so is computationally expensive, especially when considering additions, since there are many potential candidate loops. Therefore we show in the following how to construct valid loops for addition, and to determine valid loops for removal, efficiently.
@@
The addition of a loop to a polycube loop structure $\mathcal{L}$ introduces a strip of faces in the corresponding polycube $Q$, see Figure 9. We say that a loop is valid if it can be added to $\mathcal{L}$ such that $\mathcal{L}$ remains a polycube loop structure.
To effectively find valid loops for $\mathcal{L}$, we develop a directed graph structure that enables the enumeration of all possible valid loops. We say that any two loops are combinatorially equivalent if their intersection patterns (the loop segments they intersect) in $\mathcal{L}$ are equivalent. As such, we observe that all possible (valid or invalid) loops that could be added to a loop structure are characterized by the edge-graph of $\mathcal{L}$, which is defined as follows: there is a vertex for each loop segment in $\mathcal{L}$, and there is a directed edge from vertex $u$ to vertex $v$ if the two corresponding loop segments bound the same loop region in $\mathcal{L}$. We call this directed graph $G_{L}$. Note that this graph is directed since we care about the orientation of the loops. For an oriented loop, we assume that the right side (following its direction) is the positive ($+$) side, and the left side is the negative (--) side.
The validity of a loop in $G_{L}$ also depends on its label. Therefore, we will construct three different graphs, $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$, one for each possible label, where each graph is obtained by removing directed edges from $G_{L}$. The idea is that a loop with label $\Delta\in\{X,Y,Z\}$ is valid if and only if it corresponds to a simple (oriented) cycle in $G_{V}^{\Delta}$. A cycle is simple if it visits every loop region of $\mathcal{L}$ at most once.
Local constraints. When we add a loop $\lambda$ to $\mathcal{L}$, all loop regions that are visited will be split into two loop regions. All other loop regions are unaffected. Note that the splitting of a loop region, results in two new loop regions with the topology of a disk, so Condition 4 of Definition 3.2 is never violated when adding a loop.
The other local conditions of Definition 3.2 (1-3) can be guaranteed by constructing the correct graphs $G_{V}^{X}$, $G_{V}^{Y}$, and $G_{V}^{Z}$. First of all, Condition 1 can always be guaranteed by construction, as there always exists a combinatorially equivalent loop with that property. For Conditions 2 and 3, observe that every (directed) edge in $G_{L}$ corresponds to an individual split of a loop region into two loop regions. For each label $\Delta\in\{X,Y,Z\}$ of the loop we can simply directly check, for every directed edge in $G_{L}$, if the corresponding split results in valid loop regions according to Conditions 2 and 3. We include a directed edge in $G_{V}^{\Delta}$ if and only if the corresponding split is valid. There are three cases to consider based on the number of $\Delta$-loop segments within the loop region. If the loop region contains two $\Delta$-loop segments, the edge must separate the two existing loops. If the loop region contains one $\Delta$-loop segment, the edge must not intersect the existing $\Delta$-loop segment. If the loop region contains no $\Delta$-loop segments, there are no conflicts, and all edges are permitted. There is one consistent orientation of the edges if the loop region contains one or two $\Delta$-loop segments, such that the new edge does not conflict with existing edges. Otherwise, edges can be directed in either direction. These three general cases are illustrated in Figure 10, where all valid outgoing edges for a single loop segment (vertex in graph $G_{V}^{X}$) are shown.
Note that we need to require that every loop region is visited at most once; otherwise, there will be a new loop region where the same loop (the added loop) occurs twice on the boundary. Since the underlying surface $\mathcal{S}$ is orientable, the side label of the two loop segments in the direction of that loop region must also be the same, which would violate Condition 3 of Definition 3.2. Thus, a loop with label $\Delta\in\{X,Y,Z\}$ satisfies Conditions 1-4 of Definition 3.2 if and only if it corresponds to a directed cycle in $G_{V}^{\Delta}$ that visits every loop region at most once (simple).
@@
Maintaining acyclicity. The only condition that remains to be checked is Condition 5 of Definition 3.2. We show that adding a loop can never introduce a cycle in the level graphs. Hence, a loop with label $\Delta\in\{X,Y,Z\}$ is valid if and only if it corresponds to a non-trivial simple directed cycle in $G_{V}^{\Delta}$.
@@
First of all, Condition 1 of Definition 3.1 cannot be violated by the removal of a loop. For the other local Conditions 2-4, observe that, when removing a loop $\lambda$ from $\mathcal{L}$, every loop segment of $\lambda$ causes two loop regions to be merged. We can easily find these loop regions by tracing $\lambda$ through $\mathcal{L}$ and then we can simply check if the resulting loop regions satisfy Conditions 2-4.
Finally, we need to check Condition 5 of Definition 3.2 after removing $\lambda$ from $\mathcal{L}$. Note that $\lambda$ corresponds to an edge in the $X$-graph. Removing $\lambda$ does not only remove this edge, but also contracts the edge (since the zones corresponding to the endpoints are merged) into a new vertex $v$. This may introduce a cycle in the $X$-graph through $v$, but this can efficiently be checked using a simple depth first search from $v$.
Figure 12: Iterative construction of a polycube map via its dual loop structure.
@@
We first consider polycubes of genus 00, where the initial polycube loop structure corresponds to a single cube. In Figure 11(a), we show a polycube that cannot be obtained from the single cube by only adding loops. We refer to this polycube as the non-incremental polycube. It is easily verified that none of the loops of the non-incremental polycube are valid loops for removal. It follows that it is impossible to construct the non-incremental polycube by iteratively adding valid loops. However, we can construct the non-incremental polycube by both adding and removing loops. In fact, we do not know if all polycubes of genus 00 can be obtained by adding and removing loops, starting from the single cube. We leave this as an open problem.
We now move to polycubes of higher genus. Adding or removing loops cannot change the orientation of a hole in the initial polycube (see Figure 11(b, c)).